# LoopGen: Training-Free Loopable Music Generation

### v3

@@ 1 Introduction
In this paper, we introduce a loop-aware generation framework that modifies the iterative inference of a non-autoregressive (NAR) model to produce seamless loops. Concretely, we adopt a circular padding strategy, replicating partial portions of the loop at both ends of the generation window, so that the model attends to the loop’s beginning while generating its ending (Figure 1). This ensures a smooth endpoint-to-onset transition, effectively creating “bridging tokens” that align the tail of the sample with its onset. Our method can be used in two ways: (1) to generate entire loopable segments from scratch, or (2) to refine the end of an existing audio sample so that it loops seamlessly. Additionally, we implement a simple beat-aware technique that constrains the total length of the loop to align with musical bars, further promoting coherent repetition.
@@ 2 Related work
Another wave of expressive and accurate models has come with the advent of diffusion models [13] such as AudioLDM [4] and [5, 6, 14, 15]. Their application has also reached audio and music and, in this, they are giving high quality results on-par with the transformer models.
@@
Loopable music remains comparatively underexplored. LoopNet [20] specifically targets the generation of seamless music loops, but it uses a feed-forward Wave-U-Net–based model tied to a limited loop dataset with high-level musical parameters. Although it provides intuitive user controls, it falls short of the general-purpose audio and stylistic diversity achieved by large-scale transformer approaches. Other work has focused on symbolic loops in MIDI: for instance, [21, 22] propose architectures that ensure MIDI segments are musically consistent when repeated. However, these methods are intrinsically different from our focus on raw audio tokens; MIDI loops require explicit pitch and instrument representations, which do not transfer to audio-generation tasks.
Finally, loopable media generation is being tackled in computer vision with tiling techniques. Models like TileGAN [23] and [24, 25] synthesize textures or images that repeat edge-to-edge without visible seams. While these visual tiling approaches share the overarching idea of boundary alignment, they operate in a spatial domain and do not directly address audio continuity or musical structure.
@@ 4.4 Signature-aware length control
A well-formed loop often sounds most musical when it aligns with full bars (e.g., 2 or 4 bars of consistent tempo). Generating loops of arbitrary length $L$ may create awkward breaks if, for instance, the tempo does not fit integer bar divisions.
@@
  • The median number of beats per bar, $\lambda$.
Then we compute a candidate bar length $l$, which we repeatedly double or halve until it fits within user-specified bounds $[\alpha,\beta]$, and finally, we run the tiled-generation procedure with this chosen $c$.
@@
$\lambda\leftarrow$ median #beats between $D$s $\triangleright$ bar of the clip
$l\leftarrow n\lambda\delta$triangleright duration of $n$ bars
@@
if $l\in[\frac{n\lambda\delta}{4},4n\lambda\delta]$ then return $l$triangleright Return if sufficiently close
elsereturn $\varnothing$ $\triangleright$ Otherwise abort (try another $C$)
@@ 5 Evaluation Metrics
When assessing looped music, a clip that sounds fine in a single pass may still have an abrupt transition when it repeats. This effect is usually not captured by standard metrics in the audio literature, such as FAD [27], which instead measures the overall semantic and technical quality of a sample.
@@ 5.1.1 Cross-entropy and perplexity
First, recall the cross-entropy measure for a sequence $X=(x_{1},x_{2},\ldots,x_{T})$. A well-trained model $\mathcal{M}$ assigns a probability $\mathcal{M}(x_{i})$ to each token $x_{i}$. The average cross-entropy $H(X)$ is:
@@
Intuitively, if $\mathcal{M}$ assigns higher probability to each token, the negative log-probability (and hence cross-entropy) will be smaller, indicating better alignment between model and data. From cross-entropy, we derive the perplexity $\mathcal{P}(X)$, a standard measure of how well a model predicts a sequence:
@@ 6.4 Human evaluation
In total, we collect $506$ data points $s$ from $18$ listening sessions. We then mean-center the scores by removing user bias via
$s^{\prime}_{u}=s_{u}-\mu_{u}+\mu_{\textrm{global}},$ (6)
where $\mu_{\textrm{global}}$ is the global rating mean and $\mu_{u}$ is the mean user’s rating. However, this adjustment has minimal impact on the overall mean ratings for both models: indicating that ratings are stable between users. Through the raw ratings, computing each user’s average rating of each model; we run a paired t-test such that

### v4

@@ 1 Introduction
Diagram showing MAGNeT’s input divided into three sections: a central tile, a left padding, and a right padding. The central tile contains the main audio tokens to be generated. The left padding duplicates the tile’s rightmost tokens, while the right padding duplicates its leftmost tokens. This structure encourages smoother looping by exposing the model to boundary conditions during inference.
@@
In this paper, we introduce a loop-aware generation framework that modifies the iterative inference of a non-autoregressive (NAR) model to produce seamless loops. Concretely, we adopt a circular padding strategy, replicating partial portions of the loop at both ends of the generation window, so that the model attends to the loop’s beginning while generating its ending (Figure 1). This ensures a smooth endpoint-to-onset transition, effectively creating “bridging tokens” that align the tail of the sample with its onset. Our method can be used in two ways: (1) to generate entire loopable segments from scratch, or (2) to refine the end of an existing audio sample so that it loops seamlessly. Additionally, we implement a beat-aware technique that constrains the total length of the loop to align with musical bars, further promoting coherent repetition.
@@ 2 Related work
Another wave of expressive and accurate models has come with the advent of diffusion models [13] such as AudioLDM [4] and [5, 6, 14, 15, 16]. Their application has also reached audio and music and, in this, they are giving high quality results on-par with the transformer models.
@@
Loopable music remains comparatively underexplored. LoopNet [21] specifically targets the generation of seamless music loops, but it is tied to a limited dataset of loops which falls short of the general-purpose of free-form approaches. Other work has focused on symbolic loops in MIDI [22, 23], proposing architectures that ensures segments are musically consistent when repeated. However, these methods are intrinsically different from raw audio tokens; MIDI loops require explicit pitch and instrument representations, which do not transfer to audio-generation tasks. Recently, DITTO [24] introduced an inference-time optimization that allows fine-grained control, including looping, over text-to-music diffusion models. While DITTO is notable for its high output quality, it requires memory comparable to a full fine-tuning, and it slows down inference by a factor of 100$\sim$ 100$\times$ 100 ×.
Finally, loopable media generation is being tackled in computer vision with tiling techniques. Models like TileGAN [25] and [26, 27] synthesize textures or images that repeat edge-to-edge without seams. While these visual approaches share the overarching idea of boundary alignment, they do not directly address audio continuity or musical structure.
@@ 4.2 Inference algorithm
Schematic showing multiple steps of the inference algorithm. At each iteration, the central tile generates tokens, which are copied into the left or right padding in the next step. Over time, the full sequence is filled, and both padding areas end up mirroring the edges of the central tile. This iterative process reinforces loop consistency.
@@ 4.4 Signature-aware length control
A well-formed loop often sounds most musical when it aligns with full bars (e.g., 2 or 4 bars of consistent tempo). Generating loops of arbitrary length may create awkward breaks if, for instance, the tempo does not fit integer bar divisions.
@@
  • The median number of beats per bar, $\mu$.
We use these to compute the duration of a bar that the prompt $C$ implies. We want the duration of the entire loop $l$ to be a) an exact multiple (or submultiple) of the duration of a bar; b) constrained in a time interval $[\alpha,\beta]$. To achieve this, we repeatedly double or halve the initial candidate length $l$ until it fits these constraints.
@@
$\mu\leftarrow$ median #beats between $D$s $\triangleright$ bar of the clip
$l\leftarrow n\mu\delta$triangleright duration of $n$ bars
@@
if $l\in[\frac{n\mu\delta}{4},4n\mu\delta]$ then $\triangleright$ Too far from $n$ bars?
     return $l$triangleright Return if sufficiently close
else
     return $\varnothing$ $\triangleright$ Otherwise abort (try another $C$)
@@ 5 Evaluation Metrics
When assessing looped music, a clip that sounds fine in a single pass may still have an abrupt transition when it repeats. Standard metrics such as FAD [29], which measure the overall distributional similarity between generated and real audio using neural embeddings, may miss such artifacts. Because FAD operates over full-length clips, it can overlook localized issues such as discontinuities at the loop boundary. To address this, we propose a metric that focuses directly on the transition seam.
@@ 5.1 Seam perplexity
Line plot comparing cross-entropy of audio tokens around a loop point for several generation methods. Most methods show a sharp spike in cross-entropy at the loop seam, indicating poor loop quality, before quickly decreasing. LoopGen, the proposed method, avoids this spike entirely.
@@ 5.1.1 Cross-entropy and perplexity
First recall that, for a sequence $X=(x_{1},x_{2},\ldots,x_{T})$, the average cross-entropy $H(X)$ is:
@@
Intuitively, if $\mathcal{M}$ assigns higher probability to each token, the cross-entropy will be smaller, indicating better alignment between model and data. From cross-entropy, we derive the perplexity $\mathcal{P}(X)$, a standard measure of how well a model predicts a sequence:
@@ 6.3 Final results
Ridgeplot showing the distribution of seam perplexity scores for several variants. The plot visualizes the data from Table 2. MAGNeT Tiled shows slightly lower perplexity than LoopGen, though their distributions are similar. In contrast, baseline methods show much higher perplexity, indicating worse performance at loop transitions.
@@ 6.4 Human evaluation
Histogram comparing the distribution of listening test rankings for our technique versus the baseline. Our method shows a strong skew toward higher rankings, with its mean near rank 4. The baseline is skewed toward lower rankings, averaging around rank 2.
@@
In total, we collect $506$ data points $s$ from $18$ listening sessions. Computing each user’s average rating of each model; we run a paired t-test such that
@@ 7 Conclusions
8 Acknowledgments
This work is supported by Sapienza University of Rome via the Seed of ERC grant “MINT.AI”, cup B83C25001040001. Furthermore, we thank all of the participants in the human evaluation test.
@@ Appendix A Seam Perplexity’s Error Margins
Ridgeplot showing the distribution of seam perplexity scores for various audio generation methods on 10-second samples (compared to 30-second samples in the main paper). The data corresponds to Table 3. MAGNeT Tiled again shows slightly better perplexity than LoopGen, while all baseline methods have much higher perplexity, indicating poorer loop quality.
@@ Appendix C Beat alignment algorithm visualization
Diagram showing a waveform with clearly visible peaks representing beats. The first few beats are labeled as if detected by a beat detection algorithm. Based on this beat timing, the beat alignment algorithm determines the optimal point to cut the audio sample, ensuring musical continuity.